bill has 25 red tiles, 10 green tiles, and 15 blue tiles. he selects a tile at random, returns it, and selects a second tile. what is the probability that the first tile is blue and the second tile is green?
Answer by stanbon(75887)

bill has 25 red tiles, 10 green tiles, and 15 blue tiles. he selects a tile at random, returns it, and selects a second tile. what is the probability that the first tile is blue and the second tile is green?
-------
P(blue then green) = (15/25)(10/25) = 150/225 = 2/3
===================
Cheers,
Stan H.
===================